OpenStudy (djbreezy):

What is the solution to the system of equations? 4X+Y=2 X-Y=3 ( , ) @black_ninja123

OpenStudy (anonymous):

If X-Y=3..how can you change this equation to show what Y equals?

OpenStudy (djbreezy):

idk this..... @Black_ninja123

OpenStudy (black_ninja123):

I'm not good at this

OpenStudy (anonymous):

IF X -Y =3....then if we add Y to both sides we get X= 3 +Y.....now subtract 3 from both sides and what do you get?

OpenStudy (djbreezy):

X=6+y

OpenStudy (djbreezy):

@nelsonjedi

OpenStudy (anonymous):

If you subtract 3 from both sides of x=3+y ????

OpenStudy (djbreezy):

I really have a hard time understanding this...

OpenStudy (anonymous):